OpenStudy (anonymous):

What equation shows this equation in slope intercept form? y= -4/7x+9

OpenStudy (anonymous):

That's already in slope intercept form. Remember, slope intercept form is the equation of the form: y = mx + b. b is the y intercept, and m the slope, hence the name "slope intercept". In this case, m = -4/7 and b = 9